A pharmaceutical company receives large shipments of aspirin tablets. The acceptance sampling plan is to randomly select and test 26 ​tablets, then accept the whole batch if there is only one or none that​ doesn't meet the required specifications. If a particular shipment of thousands of aspirin tablets actually has a 2​% rate of​ defects, what is the probability that this whole shipment will be​ accepted? The probability that this whole shipment will be accepted is . ​(Round to three decimal places as​ needed.)
Answer #1

p = 0.02

n = 26

P(whole shipment will be accepted) = P(X < 1)

                                                        = P(X = 0) + P(X = 1)

                                                        = 26C0 * 0.020 * 0.9826 + 26C1 * 0.021 * 0.9825

                                                        = 0.905 (ans)